( * ) If (sin a + cosec a) square + (cos a + sec a) square = k + tan square a + cot square a then k =

( * )
If (sin a + cosec a) square + (cos a + sec a) square = k + tan square a + cot square a then k =

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

(sin a + cosec a)2 + (cos a + sec a)2 = k + tan2 a + cot 2 a
sin2a + cosec2a + 2 + cos2 a + sec2 a + 2 = k + tan2 a + cot 2 a
3 + cosec2a – cot 2 a +sec2 a - tan2 a = k
=> k =5